# Quantum accessible information and classical entropy inequalities

### v2

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed and proved. Via the optimality criterion, this provides also the first proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
@@ Introduction
The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, and important in particular for applications in quantum cryptography, see e.g. the survey [14]. The problem was solved in some special (notably, symmetric) cases but in general remains open. In section (2) of this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information. It turns out that the optimality criterion when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality, as explained in section 3.
@@
In the subsection 4.5 we derive and prove the tight entropy inequality (75) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids. The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one.
The proofs of the inequality (27), given in subsection 5.1, as well as of the inequalities (66), (72) in subsection 5.2 was obtained by the second author as a response to the hypotheses put forward by the first author. Via the optimality criterion of section 2, it also implies the validity of the conjectures concerning accessible information for quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2, corollary 2 to theorem 3).
@@
Let $\mathcal{H}$ be a finite dimensional Hilbert space, $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ the convex set of quantum states (density operators on $\mathcal{H}$).  Ensemble of quantum states is a collection =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ where $\rho_{j}\in\mathfrak{S,}$ and $\pi=\left\{\pi_{j}\right\}$ is a probability distribution. Let =$\mathcal{M}=\left\{M_{k}\right\}_{k=1,\dots,n}$ be an observable i.e. a collection of operators on $\mathcal{H}$ such that $M_{k}\geq 0,\,\sum_{k=1}^{n}M_{k}=I$ (unit operator). The joint distribution of “input” $j$ and “output” $k$ is then
@@
Introducing the dual ensemble =$\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
@@
$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\mathrm{\,}\,\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,$ (7)
@@
$-\sum_{j}\left\langle\phi_{k}^{0}|M_{j}^{\prime}|\phi_{k}^{0}\right\rangle\log\left\langle\phi_{k}^{0}|M_{j}^{\prime}|\phi_{k}^{0}\right\rangle=\left\langle\phi_{k}^{0}|\Lambda_{0}|\phi_{k}^{0}\right\rangle,\quad k=1,\dots.n.$ (8)
@@
$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,\quad\psi\in\mathcal{H},$
@@
$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle$
@@
$-\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (16)
@@
    for a problem which has a hypothetical solution, the conditions can be used to verify its actual optimality. One inserts the conjectured $p_{k}^{0},\sigma_{k}^{0}$ into (11) to compute Hermitean $\Lambda_{0}$. Then the main difficulty will be proof of the entropy inequality (7). In the continuous variable version, this was our strategy in the solution of the related Gaussian optimizers conjecture for the capacity of Gaussian measurement channels [8], with the new generalizations of the famous log-Sobolev inequality as the entropy inequality.
@@ 3 Example: two pure states
[…]
@@
is attained by the optimal observable =$\mathcal{M=}\left\{M_{+},M_{-}\right\}$ where $M_{\pm}=\left|e_{\pm}\rangle\langle e_{\pm}\right|$ and $|e_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|e_{0}\rangle\pm|e_{1}\rangle\right)$ is an orthonormal basis [^3] Strictly speaking, the paper [10] contains the check of local optimality.. Here
@@
$K(\sigma)=-\sum_{\pm}\left|e_{\pm}\rangle\langle e_{\pm}\right|\log\left\langle e_{\pm}|\sigma|e_{\pm}\right\rangle$
@@
The inequality (23) becomes equality for $t=p$ and $t=1-p.$
@@
Thus $h(t)-g(t,p)=0$ for $t=p,$ and $\frac{d}{dt}\left[h(t)-g(t,p)\right]$ has the same sign as $t-p.$ Hence $h(t)-g(t,p)\geq 0.$
Theorem 1 then implies the global optimality of the observable =$\mathcal{M=}\left\{M_{+},M_{-}\right\}.$ Also
@@ 4.1 The lower bound for the Shannon entropy: formulation and discussion
$-\sum_{j=1}^{m}t_{j}\log t_{j}\geq\mu_{0}(p)\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\mu_{1}(p),$ (27)
@@
The inequality (27) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection. The proof of (27) will be given in section 5.1. Note that in the case $m=2$ the inequality (27) reduces to (23).
@@
$-\sum_{j=1}^{m}t_{j}\log t_{j}\geq\log m-\frac{m\log m}{2\left(m-1\right)}D(P;P_{U}),$ (32)
@@ 4.2 The conjecture: “quantum pyramids”
We will consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
@@
In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2]. In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
$K\left(\sigma\right)=-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log\mathrm{\,}\,\langle e_{j}|\sigma|e_{j}\rangle.$
From (41), (39) we obtain the dual ensemble =$\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
@@
-- $0\displaystyle 2\frac{1+\left(m-1\right)s}{m}\ln\left[1+\left(m-1\right)s\right]-2\frac{m-1}{m}\left(1-s\right)\ln\left|1-s\right|=$ 0 .
@@
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (57)
Since $\mu_{0}(p)\geq 0,$ this is equivalent to (27) where $t_{j}=\left|z_{j}\right|^{2}$
@@
Taking into account theorem 2, the first statement was in fact proved above in this section. For accessible information we use the relation (58). A routine calculation using (52), (37), (56) gives (45) which is the first line in (59) for moderately acute pyramids. For strongly acute pyramid, using (55), one gets
@@ 4.4 The case of obtuse pyramids
When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different. To describe it let us first consider the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf. [5]. Then the optimal observable is =$\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [12]. In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
$-\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (60)
@@
Most interesting is the case of “nearly flat” trine when $06140<r_{0}<\bar{r}_{0}\approx 0,$ 0614 . Then, as shown by P. Shor in [13], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (33) with $r_{0}<r_{1}$ $(b<0)$ The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
@@
For complex $z_{j}$ , $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,\,$
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (66)
@@
Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$. In (66) the parameter $p$ varies in the range $[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below. The accessible information in this case is given by the same expression (45) as in the case of moderately acute pyramids.
@@
$p\equiv a^{2}<\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}\equiv p(m).$ (71)
@@
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (72)
The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
@@
The proof of the inequalities is given in section 5.2.

### v3

@@ Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed. Via the optimality criterion, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
@@ Introduction
The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, see e.g. [6], [16], [9], [11], and important in particular for applications in quantum cryptography. The problem was solved in some special (notably, symmetric) cases but in general remains open. In section (2) of this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information. It turns out that the optimality criterion when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality, as explained in section 3.
@@
In the subsection 4.5 we derive the tight entropy inequality (82) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids. The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and tentatively is the optimal one.
The proposed entropy inequalities may be more convenient for numerical check than the direct optimization, however proving them analytically is a challenge. In section 5, the suggestions for proofs proposed by the second author as a response to the hypotheses (27), (67), (79) put forward by the first author, are presented. Via the optimality criterion of section 2, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
@@
Let $\mathcal{H}$ be a finite dimensional Hilbert space, $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ the convex set of quantum states (density operators on $\mathcal{H}$).  Ensemble of quantum states is a collection $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ where $\rho_{j}\in\mathfrak{S,}$ and $\pi=\left\{\pi_{j}\right\}$ is a probability distribution. Let $\mathcal{M}=\left\{M_{k}\right\}_{k=1,\dots,n}$ be an observable i.e. a collection of operators on $\mathcal{H}$ such that $M_{k}\geq 0,\,\sum_{k=1}^{n}M_{k}=I$ (unit operator). The joint distribution of “input” $j$ and “output” $k$ is then
@@
Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
@@
[…]
@@
[…]
@@
$\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,\quad\psi\in\mathcal{H},$
@@
[…]
@@
$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (16)
@@
    for a problem which has a hypothetical solution, the conditions can be used to verify its actual optimality. One inserts the conjectured $p_{k}^{0},\sigma_{k}^{0}$ into (11) to compute Hermitean $\Lambda_{0}\$. Then the main difficulty will be proof of the entropy inequality (7). In the continuous variable version, this was our strategy in the solution of the related Gaussian optimizers conjecture for the capacity of Gaussian measurement channels [8], with the new generalizations of the famous log-Sobolev inequality as the entropy inequality.
@@ 3 Example: two pure states
The case of ensemble of two states was treated previously in [12], [10]. For simplicity of demonstration we take two pure equiprobable states. Let $|e_{0}\rangle,|e_{1}\rangle$ be an orthonormal basis in two-dimensional Hilbert space. Consider ensemble $\mathcal{E}=\left\{\pi_{\pm},\rho_{\pm}\right\},$ where $\pi_{\pm}=\frac{1}{2},\rho_{\pm}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|$ with $|\psi_{\pm}\rangle=\cos\alpha/2|e_{0}\rangle\pm\sin\alpha/2|e_{1}\rangle,\,0\leq\alpha\leq\pi.$ As argued in [12] the accesible information
@@
is attained by the optimal observable $\mathcal{M=}\left\{M_{+},M_{-}\right\}$ where $M_{\pm}=\left|e_{\pm}\rangle\langle e_{\pm}\right|$ and $|e_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|e_{0}\rangle\pm|e_{1}\rangle\right)$ is an orthonormal basis [^3] Strictly speaking, the paper [12] contains the check of local optimality.. Here
@@
[…]
@@
The inequality (23) becomes equality for $t=p$ and t=1p.t=1-p.
@@
Thus h(t)g(t,p)=0h(t)-g(t,p)=0 for $t=p,$ and $\frac{d}{dt}\left[h(t)-g(t,p)\right]$ has the same sign as tp.t-p. Hence $h(t)-g(t,p)\geq 0.$
Theorem 1 then implies the global optimality of the observable $\mathcal{M=}\left\{M_{+},M_{-}\right\}.$ Also
@@ 4.1 The lower bound for the Shannon entropy: formulation and discussion
$\sum_{j=1}^{m}t_{j}\log t_{j}\geq\mu_{0}(p)\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\mu_{1}(p),$ (27)
@@
The inequality (27) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection. The suggestion for a proof of (27) will be given in section 5.1. Note that in the case $m=2$ the inequality (27) reduces to (23).
@@
$\sum_{j=1}^{m}t_{j}\log t_{j}\geq\log m-\frac{m\log m}{2\left(m-1\right)}D(P;P_{U}),$ (32)
@@ 4.2 The conjecture: “quantum pyramids”
We will consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
@@
In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2]. In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form $\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
@@ 4.3 Checking the optimality conditions: the case of acute pyramids
[…]
From (41), (39) we obtain the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
@@
- $2\frac{1+\left(m-1\right)s}{m}\ln\left[1+\left(m-1\right)s\right]-2\frac{m-1}{m}\left(1-s\right)\ln\left|1-s\right|=0.$
@@
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (57)
Since $\mu_{0}(p)\geq 0,$ this is equivalent to (27) where $t_{j}=\left|z_{j}\right|^{2}$ .
@@
For accessible information we use the relation (58). A routine calculation using (52), (37), (56) gives (45) which is the first line in (59) for moderately acute pyramids. For strongly acute pyramid, using (55), one gets
@@ 4.4 The case of obtuse pyramids
When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different. To describe it let us first consider the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf. [5]. Then the optimal observable is $\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [14]. In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
$\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (60)
@@
Most interesting is the case of “nearly flat” trine when $0<r_{0}<\bar{r}_{0}\approx 0,0614.$ Then, as shown by P. Shor in [15], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (33) with $r_{0}<r_{1}$ $(b<0)$ . The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
@@
For complex $z_{j}$ satisfying
$\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ (66)
it holds
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (67)
@@
Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$. However, as the following lemma shows, we can always restrict to the case of real $z_{j}$.
Lemma 1.
Let $\tilde{\mu}_{0},\tilde{\mu}_{1}$ be real numbers. If the inequality
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}\leq\tilde{\mu}_{1}$ (70)
holds for all real $z_{j}$ satisfying the condition (66), then it holds for all complex $z_{j}$ satisfying the same condition.
If it holds for all real $z_{j}$ satisfying the condition (66) and
$\mathit{\sum_{j=1}^{m}z_{j}=0,}$ (71)
then it holds for all complex $z_{j}$ satisfying the conditions (66), (71).
Proof.
Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (66) amounts to
$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad$ (72)
and (71) amounts to
$\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (73)
Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, hence
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
$+$ $P\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}+(1-P)\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}$
$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}\right]$
$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}\right]$
$\leq$ $\tilde{\mu}_{1},$
because
$\sum_{j=1}^{m}\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}=\sum_{j=1}^{m}\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}=1,\quad$
and by assumption (70) holds for real $z_{j}$ satisfying (66). If additionally the complex $z_{j}$ satisfy (71) then
$\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
and the second statement of the lemma follows. ∎
In (67) the parameter $p$ varies in the range $\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below. The accessible information in this case is given by the same expression (45) as in the case of moderately acute pyramids. For future use we note that
$\tilde{\mu}_{1}\left(\frac{m-1}{m}\right)=\log m-\frac{m-2}{m}\log(m-1).$ (74)
@@
$p\equiv a^{2}$ $<$ $\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}$ (77)
$=$ $\frac{(m-1+\tau_{o}(m))^{2}}{m(m-1+\tau_{o}(m)^{2})}\equiv p(m).$
One can check that equation (76) with $\tau=\tau_{o}(m)$ is equivalent to
$\tilde{\mu}_{1}(p(m))=1.$ (78)
@@
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (79)
Due to lemma 1 one can restrict here to real $z_{j}$ , $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1.$ The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
@@
The suggestion for a proof of the inequalities is given in section 5.2.